OpenStudy (anonymous):

Find the common ratio for the following sequence. 27, 9, 3, 1, ... -1?

OpenStudy (anonymous):

it is Gp

OpenStudy (anonymous):

R = second term/first term

OpenStudy (anonymous):

so r = 9/27 = 1/3

OpenStudy (anonymous):

so 1/3?

OpenStudy (jiteshmeghwal9):

\[(1-x)(1+x+x^2.....+x^n)=1-x^{n+1}\]where n=0,1,2...